# On the Eligibility of LLMs for Counterfactual Reasoning: A Decompositional Study

### v1

@@ 1 Introduction
Large language models (LLMs) have exhibited remarkable proficiency across a diverse range of tasks, including natural language understanding devlin2019bert ; 10.1145/3711680 and multimodal reasoning hu2017learning ; lu2022learn ; yang2023mm . Despite these advancements, concerns persist regarding their reasoning and generalization capabilities. A particularly challenging aspect of model evaluation is Counterfactual Reasoning, i.e., the ability to adjust responses when presented with modified premises pearl2018book (e.g., What is the outcome in a hypothetical condition?). Investigating the counterfactual reasoning of LLMs provides an interpretable step to understand their adaptability under hypothetical alterations to input conditions gatfaithful ; huang-etal-2024-clomo .
Prior studies have demonstrated that LLMs often struggle with counterfactual reasoning and frequently fail to maintain logical consistency or adjust to context shifts li-etal-2023-counterfactual ; nguyen-etal-2024-llms ; wang-etal-2024-survey . While these works highlight notable performance gaps, they lack a standardized framework for systematically analyzing and understanding counterfactual behaviors in LLMs. Consequently, it remains unclear what factors most significantly impact LLM performance in counterfactual scenarios. Furthermore, counterfactual reasoning has often been evaluated in a direct and monolithic manner, primarily by introducing interventions and assessing model responses li-etal-2023-counterfactual , without grounding the analysis in the underlying causal structure that gives rise to such interventions. This overlooks the foundational role of causal modeling. Specifically, the identification of causal variables and their dependencies are essential for understanding counterfactuals.
@@
To address these gaps, we outline a Decompositional Strategy that breaks down the analysis of counterfactual reasoning. Our approach departs from prior work that focuses solely on counterfactual generation. Instead, we begin by examining the causal structure of factual conditions, which serves as the necessary foundation for valid counterfactual reasoning.
@@
To support our decompositional study, we construct a benchmark by collecting and curating 11 counterfactual datasets across diverse tasks, including natural language understanding, mathematics, programming, and vision-language reasoning. We curate each dataset by extracting factual and counterfactual variables, identifying causal elements, and constructing corresponding causal graphs as reference structures for evaluation purpose. In experiments, we test the performance of leading LLMs across each decompositional stage to analyze their sufficiency in handling individual reasoning components. Based on the observed performance among these decomposed evaluations, we propose targeted improvements, such as integrating modality-specific function-calling interfaces within a tool-augmented learning paradigm, to address critical reasoning bottlenecks. Additionally, we evaluate the impact of different elicitation (prompting) strategies, including Chain-of-Thought (CoT) wei2022chain , Chain-of-Thought with Self-Consistency (CoT-SC) wang2022self , and Tree-of-Thought (ToT) 10.5555/3666122.3666639 reasoning. Collectively, our evaluations provide a solid step for understanding and enhancing LLMs in complex reasoning tasks and imaginative scenarios.
@@ 2 Related Work
Counterfactual Reasoning. A fundamental component of causal inference is to examine hypothetical scenarios, which addresses the question: What would have occurred had a particular factor or decision differed? pearl2018book This method facilitates causal analysis by comparing observed outcomes with those projected under alternative conditions. Empirical research has demonstrated the broad applicability of counterfactual reasoning across multiple domains, including healthcare, business, and fairness gvozdenovic2021causal ; kyrimi2025counterfactual ; gow2016causal ; kasirzadeh2021use ; koonce2011judging . This stands in contrast to many contemporary AI systems, which predominantly rely on statistical correlations while lacking robust capacities for abstract reasoning and causal inference jiao2024causal .
@@
Counterfactual Reasoning in AI and NLP. Counterfactual reasoning has emerged as a powerful framework for enhancing model interpretability and causal understanding in AI and NLP. In medical AI, SyncTwin NEURIPS2021_19485224 proposed a counterfactual estimation framework that constructs synthetic patient data to predict potential outcomes under alternative treatments. In NLP, a Counterfactual Reasoning Model (CRM) feng-etal-2021-empowering is developed using LLMs to generate contrastive samples, improving sentiment analysis and inference tasks. There are also order-faithfulness metrics gatfaithful to evaluate causal explanations in black-box models. These contributions demonstrate the versatility of counterfactual methods in improving model transparency and reliability across domains.
@@
LLMs and Elicitation. Recent advances have significantly enhanced LLMs’ reasoning capabilities through several elicitation (prompting) approaches. The introduction of Chain-of-Thought (CoT) prompting wei2022chain and its extensions, the Self-Consistency CoT wang2022self and Tree-of-Thought (ToT)  10.5555/3666122.3666639 , have enabled more structured and reliable multi-step reasoning. Complementing these prompting techniques, reinforcement learning approaches using reward models lightman2023let have further improved reasoning quality through fine-grained supervision. These innovations collectively represent a paradigm shift from simple pattern recognition to deliberate, verifiable reasoning in LLMs.
@@ 3.1 Preliminary: From Causality to Counterfactual Reasoning
Causality. Causality depicts the dependencies about how one variable influences another, i.e., the underlying causal effects. There are four types of variables commonly used in causal analysis: exposure, covariate, mediator, and outcome. Specifically: (1) Exposure (or treatment, intervention, denoted $X$) refers to the action or condition imposed on a system; (2) Outcome ($Y$) denotes the resulting response or effect influenced by the exposure; (3) Covariate ($Z$) are pre-treatment variables that may influence both $X$ and $Y$; (4) Mediator ($M$) lies on the causal pathway from $X$ to $Y$, representing intermediate mechanisms through which the exposure exerts its influence.
@@
Causal Graph. The relationships among exposure, covariate(s), mediator(s), and outcome(s) can be formally represented using a directed acyclic graph (DAG), commonly named a causal graph that captures causal relationships, where the exposure $X$ influences the outcome $Y$ both directly and indirectly through a mediator $M$. Covariate $Z$ may affect the relationships among $X$, $M$, and $Y$.
@@
Given an observed data instance $(X=x,Z=z,M=m,Y=y)$, what would the outcome $Y$ have been if the exposure $X$ were set to a different value $x^{\prime}$, while holding the covariate $Z$ fixed?
@@
[…]
@@ 3.2 Decompositional Evaluation Task
We define four evaluation tasks to systematically analyze LLMs’ counterfactual reasoning ability:
@@
Our evaluation spans these decompositional tasks across a variety of modalities (e.g., text, images, mathematical symbols, codes) and LLMs (e.g., GPT, Claude, Llama). Our ultimate goal is to quantify performance variations under different conditions and pinpoint how LLMs succeed or struggle in counterfactual reasoning.
@@ 3.3 Benchmarking Counterfactuals
Next, we introduce the datasets we leverage and the methodology we adopt to curate them for benchmarking LLMs’ counterfactual reasoning capabilities.
@@
[…]
Data Sources and Use Cases. As shown in Table 1, we collect a diverse set of datasets to ensure broad coverage across various NLP tasks and modalities. The included use cases are: (1) Question Answering, evaluated using CRASS frohberg2021crass , CVQA-Bool zhang2024if , and MalAlgoQA sonkar2024malalgoqa , which involve answering general-purpose textual or visually grounded questions; (2) Text Parsing, using CLOMO huang2023clomo for logical structure reconstruction and RNN-Typolog ravfogel2019studying for syntactic structure understanding; (3) Reasoning Tasks, with CVQA-count zhang2024if for numerical reasoning and Arithmetic wu2024reasoning for symbolic arithmetic computation; (4) Multimodal Matching, represented by the COCO dataset le2023coco for image-text alignment; (5) Code-based Tasks, including HumanEval-Exe chen2021evaluating for execution simulation, Open-Critic vezora_code_preference_pairs_2024 for generation, and Code-Preference vezora_code_preference_pairs_2024 for summarization.
@@
Our Curation. To adapt these datasets for counterfactual reasoning evaluation, we conduct a careful manual curation process to augment each instance with three additional aspects of information relevant to Tasks I–III (Section 3.2). Specifically, we begin by identifying and annotating the causal variables $(X,Z,M,Y)$ from the original data, questions, or descriptions. Using these annotations, we construct a DAG to represent the underlying causal structure of each data instance.
All selected datasets naturally support counterfactual interventions (e.g., “What if…?” queries or hypothetical scenarios). For each instance, we parse and extract the intervened variables and, guided by the previously constructed DAG, annotate the corresponding counterfactual outcomes and construct a matched counterfactual graph. We provide the curated instances in Appendix A.
This curation ensures that each benchmark instance is enriched with explicit causal and counterfactual structures. During the evaluation, we not only test LLMs’ overall counterfactual reasoning performance but also decompose the evaluation into Tasks I–III, leveraging our curated causal variables, DAGs, counterfactual interventions, and corresponding outcome graphs.
@@ 4 Experiment
We aim to empirically answer two research questions: RQ1: How well do LLMs perform when their counterfactual reasoning is decomposed into distinct reasoning tasks? RQ2: What auxiliary techniques (e.g., calling tools, prompting strategies) can improve LLMs’ counterfactual reasoning? We defer the detailed settings into Appendix C and discussion of limitation into Appendix D.
LLMs. Given the multimodal nature of counterfactual tasks, we evaluate several Visual LLMs, including GPT-4o openai2024gpt4o , Qwen-VL Qwen-VL , LLaMA-3.2-11B Meta2024Llama , Gemini-Pro gemini25pro , DeepSeek-VL lu2024deepseekvl .
@@ 4.1 LLM Performance on Decompositional Tasks (RQ1)
Setting. We evaluate LLMs independently on each decompositional task. For each task, we explicitly provide the ground-truth outputs from the preceding tasks to isolate and measure LLMs’ capabilities specific to that task. For example, when assessing models’ ability to construct causal graphs, we supply the original inputs along with the ground-truth causal variables, ensuring that the evaluation focuses solely on graph-building performance.
Task I: Causal Variable Identification. Table 2 presents the performance of LLMs on identifying causal variables ($X,Z,M,Y$). We observe that model performance is strongly influenced by the modality of the dataset. Specifically, datasets involving more complex modalities—such as images, mathematical symbols, or code—tend to obviously reduce LLM accuracy (e.g., $<$0.7 F1 on Open-Critic), even when variables like $X,Z,Y$ are explicitly present in the context.
@@
Table 2: LLMs’ performance (F1, scaled to 100%) in causal variable identification. We report mean±$\pm$±std across all instances for each variable, since $Z,M$ may include multiple entities.
Dataset GPT-4o Qwen-VL Llama-3.2-11B Gemini-Pro DeepSeek-VL
$v_{1}$ $v_{2}$ $v_{1}$ $v_{2}$ $v_{1}$ $v_{2}$ $v_{1}$ $v_{2}$ $v_{1}$ $v_{2}$
$v_{1}=X$ (Exposure), $v_{2}=Z$ (Covariate)
CRASS 92.5±3.2 90.8±4.1 87.3±5.6 85.4±6.2 74.8±8.3 72.5±7.9 88.6±4.5 86.2±5.3 89.5±4.8 87.1±5.0
CLOMO 89.6±4.7 87.3±5.2 83.5±6.8 81.9±7.2 70.2±9.6 68.7±8.9 84.3±5.9 82.8±6.5 86.4±5.3 84.2±5.7
RNN-Topo 87.2±5.1 85.9±5.6 80.4±7.3 78.6±7.8 68.5±10.2 66.9±9.5 81.7±6.4 80.1±6.9 83.8±5.8 82.0±6.2
CVQA-Bool 76.4±8.6 74.8±9.1 72.3±9.7 70.5±10.2 56.8±12.4 54.2±13.1 68.5±10.8 66.9±11.5 70.7±9.4 68.3±10.1
CVQA-Count 72.1±9.5 70.6±10.2 68.9±10.6 67.2±11.3 52.6±13.7 50.8±14.2 65.8±11.5 63.7±12.3 67.4±10.3 65.1±11.2
COCO 70.5±10.8 68.9±11.5 67.2±11.9 65.4±12.6 48.3±15.2 46.7±15.8 62.6±12.9 60.9±13.7 65.9±11.6 63.2±12.5
Arithmetic 82.3±7.2 80.6±7.8 75.7±8.9 73.8±9.5 61.4±11.3 59.6±12.0 76.9±8.4 74.5±9.1 78.3±7.9 76.1±8.6
MalAlgoQA 79.8±7.9 77.5±8.4 72.9±9.5 70.6±10.1 58.7±12.2 56.4±12.9 73.5±9.0 71.2±9.7 75.9±8.5 73.4±9.2
HumanEval-Exe 68.9±11.4 66.5±12.1 63.7±12.8 61.9±13.6 45.8±16.3 43.2±17.1 59.6±13.9 57.3±14.7 62.1±12.5 59.8±13.4
Open-Critic 66.2±12.3 64.1±13.0 61.8±13.5 59.7±14.2 42.4±17.1 40.6±17.8 57.3±14.6 55.9±15.3 60.4±13.7 58.1±14.5
Code-Preference 77.5±11.9 65.3±12.6 72.9±13.1 60.5±13.9 54.3±16.7 42.1±17.4 68.4±14.3 66.5±15.0 61.2±13.1 59.3±13.9
$v_{1}=M$ (Mediator), $v_{2}=Y$ (Outcome)
CRASS 78.5±8.2 86.3±5.8 72.8±9.6 79.9±7.4 62.3±12.1 70.9±9.7 74.1±9.1 81.6±6.8 76.2±8.7 83.5±6.2
CLOMO 76.3±8.9 83.6±6.3 68.9±10.5 77.4±8.1 59.7±12.9 68.1±10.3 71.2±9.8 78.9±7.5 73.5±9.3 80.7±6.8
RNN-Topo 74.8±9.3 81.7±6.7 67.1±11.1 75.8±8.5 56.9±13.6 66.3±10.8 69.3±10.5 76.4±7.9 71.5±9.8 78.9±7.2
CVQA-Bool 64.8±12.1 72.5±9.7 59.7±13.8 66.9±11.3 48.2±16.9 56.3±14.2 57.3±14.1 63.2±12.5 60.5±13.2 68.4±10.7
CVQA-Count 61.3±13.0 68.7±10.5 56.8±14.6 63.6±12.1 44.5±17.8 53.2±15.1 54.7±15.0 60.4±13.3 57.9±14.1 65.1±11.6
COCO 59.7±13.7 66.3±11.2 54.2±15.3 61.8±12.8 41.9±18.5 50.4±15.9 52.8±15.7 58.3±14.1 55.7±14.8 62.6±12.4
Arithmetic 69.5±10.7 77.4±8.3 63.6±12.5 71.9±9.8 51.9±15.4 61.2±12.5 62.8±13.1 73.5±9.2 65.7±11.3 74.9±8.7
MalAlgoQA 67.6±11.3 74.8±8.9 61.5±13.1 69.3±10.4 49.7±16.1 58.9±13.2 60.2±13.8 70.5±10.0 63.8±12.0 72.3±9.3
HumanEval-Exe 57.8±14.5 64.3±12.0 51.9±16.2 59.7±13.6 39.8±19.3 48.5±16.7 49.7±16.5 57.0±14.8 53.6±15.5 60.5±13.1
Open-Critic 55.9±15.2 62.4±12.7 49.7±16.9 57.2±14.4 37.6±8.1 45.9±17.5 47.3±17.2 54.8±15.5 51.3±16.1 58.7±13.9
Code-Preference 56.8±14.9 67.5±12.3 50.4±16.5 78.6±14.0 38.4±19.7 67.2±17.0 48.2±16.9 76.1±15.1 52.5±15.8 59.8±13.5

### v2

@@ 1 Introduction
Large language models (LLMs) have exhibited remarkable proficiency across a diverse range of tasks, including natural language understanding (Devlin et al., 2019; Kuang et al., 2025), multimodal reasoning (Hu et al., 2017; Lu et al., 2022; Yang et al., 2023; Zhou et al., 2024; Tang et al., 2025b), and domain-specific generation (Desai et al., 2026; Kalvakuntla et al., 2026; Tang et al., 2025a). Despite these advancements, concerns persist regarding their reasoning and generalization capabilities. A particularly challenging aspect of model evaluation is Counterfactual Reasoning, i.e., the ability to adjust responses when presented with modified premises (Pearl & Mackenzie, 2018) (e.g., What is the outcome in a hypothetical condition?). Investigating the counterfactual reasoning of LLMs provides an interpretable step to understand their adaptability under hypothetical alterations to input conditions (Gat et al., 2024; Huang et al., 2024).
Prior studies have demonstrated that LLMs often struggle with counterfactual reasoning and frequently fail to maintain logical consistency or adjust to context shifts (Li et al., 2023; Nguyen et al., 2024; Wang et al., 2024). While these works highlight notable performance gaps, they lack a standardized framework for systematically analyzing and understanding counterfactual behaviors in LLMs. Consequently, it remains unclear what factors most significantly impact LLM performance in counterfactual scenarios. Furthermore, counterfactual reasoning has often been evaluated in a direct and monolithic manner, primarily by introducing interventions and assessing model responses (Li et al., 2023), without grounding the analysis in the underlying causal structure that gives rise to such interventions. This overlooks the foundational role of causal modeling. Specifically, the identification of causal variables and their dependencies are essential for understanding counterfactuals.
@@
To address these gaps, we are motivated by the structural formulation of counterfactual reasoning under the Structural Causal Model (SCM) formalism (Pearl, 2009), in which counterfactual reasoning must proceed through a sequence of regularized steps including inferring latent variables from observations, modifying the SCM via intervention, and computing the updated outcome. Accordingly, we outline a Decompositional Strategy  that breaks down the analysis of counterfactual reasoning into distinct stages. Our approach departs from prior work that focuses solely on counterfactual generation. Instead, we begin by examining the causal structure of factual conditions, which serves as the necessary foundation for valid counterfactual reasoning.
@@
To support our decompositional study, we construct a benchmark by collecting and curating 11 counterfactual datasets across diverse tasks, including natural language understanding, mathematics, programming, and vision-language reasoning. We curate each dataset by extracting factual and counterfactual variables, identifying causal elements, and constructing corresponding causal graphs as reference structures for evaluation purpose. In experiments, we test the performance of leading LLMs across each decompositional stage to analyze their sufficiency in handling individual reasoning components. Based on the observed performance among these decomposed evaluations, we propose targeted improvements, such as integrating modality-specific function-calling interfaces within a tool-augmented learning paradigm, to address critical reasoning bottlenecks. Additionally, we evaluate the impact of different elicitation (prompting) strategies, including Chain-of-Thought (CoT) (Wei et al., 2022), Chain-of-Thought with Self-Consistency (CoT-SC) (Wang et al., 2022), and Tree-of-Thought (ToT) (Yao et al., 2023) reasoning. Collectively, our evaluations provide a solid step for understanding and enhancing LLMs in complex reasoning tasks and imaginative scenarios.
@@ 2 Related Work
Counterfactual Reasoning. A fundamental component of causal inference is to examine hypothetical scenarios, which addresses the question: What would have occurred had a particular factor or decision differed? (Pearl & Mackenzie, 2018) This method facilitates causal analysis by comparing observed outcomes with those projected under alternative conditions. Empirical research has demonstrated the broad applicability of counterfactual reasoning across multiple domains, including healthcare, business, and fairness (Gvozdenović et al., 2021; Kyrimi et al., 2025; Gow et al., 2016; Kasirzadeh & Smart, 2021; Koonce et al., 2011). This stands in contrast to many contemporary AI systems, which predominantly rely on statistical correlations while lacking robust capacities for abstract reasoning and causal inference (Jiao et al., 2024).
@@
Counterfactual Reasoning in AI and NLP. Counterfactual reasoning has emerged as a powerful framework for enhancing model interpretability and causal understanding in AI and NLP. In medical AI, SyncTwin (Qian et al., 2021) proposed a counterfactual estimation framework that constructs synthetic patient data to predict potential outcomes under alternative treatments. In NLP, a Counterfactual Reasoning Model (CRM) (Feng et al., 2021) is developed using LLMs to generate contrastive samples, improving sentiment analysis and inference tasks. There are also order-faithfulness metrics (Gat et al., 2024) to evaluate causal explanations in black-box models. These contributions demonstrate the versatility of counterfactual methods in improving model transparency and reliability across domains.
@@
LLMs and Elicitation. Recent advances have significantly enhanced LLMs’ reasoning capabilities through several elicitation (prompting) approaches. The introduction of Chain-of-Thought (CoT) prompting (Wei et al., 2022) and its extensions, the Self-Consistency CoT (Wang et al., 2022) and Tree-of-Thought (ToT)  (Yao et al., 2023), have enabled more structured and reliable multi-step reasoning. These innovations collectively represent a paradigm shift from simple pattern recognition to deliberate, verifiable reasoning in LLMs (Meng et al., 2025; Liu et al., 2025).
Evaluation of Counterfactual Reasoning. Evaluating counterfactual reasoning in LLMs has garnered growing attention. However, most prior works assess counterfactual through end-to-end evaluations such as contrastive counterfactuals (e.g., “What would happen if X didn’t occur?”) (Huang et al., 2023; Frohberg & Binder, 2021; Zhang et al., 2024b; Le et al., 2023). Other studies such as MalAlgoQA (Sonkar et al., 2024) introduce the concept of algorithms to assess LLMs’ ability to reason about flawed hypothetical paths. Their setup focuses on identifying distractor rationales in multiple-choice formats, revealing LLM struggles in understanding student misconceptions Other efforts like DICE (Shrivastava & Aoyagui, 2025) and CausalProbe (Chi et al., 2024) create diagnostic benchmarks to evaluate causal sensitivity or counterfactual faithfulness in static question-answering formats, without decomposing the reasoning process into interpretable modules . Similarly, synthetic datasets have been used to analyze RNN inductive biases in agreement prediction (Ravfogel et al., 2019) but do not extend to structured counterfactual inference tasks. Compared to these works, our study introduces a modular evaluation aligned with Pearl’s structural causal model (Pearl & Mackenzie, 2018). We decompose counterfactual reasoning into four interconnected sub-tasks to open up fine-grained attribution of model failures and provide a more diagnostic and interpretable assessment of LLM reasoning capabilities.
@@ 3.1 Preliminary: From Causality to Counterfactual Reasoning
Causality. Causality depicts the dependencies about how one variable influences another, i.e., the underlying causal effects. There are four types of variables commonly used in causal analysis: exposure, covariate, mediator, and outcome. Specifically: (1) Exposure (or treatment, intervention, denoted $X$) refers to the action or condition imposed on a system; (2) Outcome ($Y$) denotes the resulting response or effect influenced by the exposure; (3) Covariate ($Z$) is the pre-treatment variable that may influence both $X$ and $Y$; (4) Mediator ($M$) lies on the causal pathway from $X$ to $Y$, representing intermediate mechanisms through which the exposure exerts its influence.
@@
Causal Graph. The relationships among exposure, covariate(s), mediator(s), and outcome(s) can be formally represented using a directed acyclic graph (DAG), commonly named a causal graph (Pearl & Mackenzie, 2018) that captures causal relationships, where the exposure $X$ influences the outcome $Y$ both directly and indirectly through a mediator $M$. Covariate $Z$ may affect $X$, $M$, and $Y$, as illustrated in Figure 2.
@@
Given an observed instance $(X=x,Z=z,M=m,Y=y)$, what would the outcome $Y$ be if the exposure $X$ were set to a different value $x^{\prime}$, while keeping the covariate $Z$ fixed?
@@
Example 3 As exemplified in Figure 2-(b), a student with $z=\textsc{low-income}$ socioeconomic status did not receive tutoring $(x=0)$, studied 5 hours per week $(m=5)$, and scored 78 on the exam $(y=78)$. We now ask: What would the student’s score have been if they had received tutoring $(x^{\prime}=1)$? We compute: (i) Simulated study time: $m^{\prime}=f_{M}(x^{\prime}=1,z=\textsc{low-income})=9$ and (ii) counterfactual score: $Y_{x^{\prime}=1}=f_{Y}(x^{\prime}=1,\ m^{\prime}=9,\ z=\textsc{low-income})=85$. We conclude: The tutoring would have increased the student’s score from 78 to 85 (Figure 2-(c)).
@@ 3.2 Decompositional Evaluation Task
Counterfactual reasoning is often described as a structured chain of analysis from identifying variables to modeling causal relations, specifying interventions, and simulating outcomes (Pearl, 2009; Bareinboim et al., 2022). Recent work in ML and LLMs further emphasizes the importance of disentangling these steps to evaluate reasoning capacity (Kiciman et al., 2023; Chi et al., 2024). Motivated by the need for decomposition, we design four evaluation tasks that reflect the full pipeline of counterfactual reasoning, with each task targeting a distinct capability in counterfactual analysis.
@@ 3.3 Benchmarking Counterfactuals
Next, we introduce the datasets we leverage for the decompositional evaluations:
@@
CLOMO (Huang et al., 2023) Text logic parsing $CIRCLE \CIRCLE \CIRCLE$ New premise for textual statement Text 1,100
RNN-Typology (Ravfogel et al., 2019) Text syntax parsing $CIRCLE \CIRCLE \CIRCLE$ New syntactic structure of sentence Text 584
[…]
COCO (Le et al., 2023) Text-image matching $CIRCLE \LEFTcircle \CIRCLE$ “What if …” condition Text,Image 17,410
Arithmetic (Wu et al., 2024) Mathematical reasoning $CIRCLE \CIRCLE \CIRCLE$ Change number base Symbol 6,000
@@
Data Sources and Use Cases. As shown in Table 1, we collect a diverse set of datasets to ensure broad coverage across various NLP tasks and modalities. The included use cases are: (1) Question Answering, evaluated using CRASS (Frohberg & Binder, 2021), CVQA-Bool (Zhang et al., 2024b), and MalAlgoQA (Sonkar et al., 2024), which involve answering general-purpose textual or visually grounded questions; (2) Text Parsing, using CLOMO (Huang et al., 2023) for logical structure reconstruction and RNN-Typolog (Ravfogel et al., 2019) for syntactic structure understanding; (3) Reasoning Tasks, with CVQA-Count (Zhang et al., 2024b) for numerical reasoning and Arithmetic (Wu et al., 2024) for symbolic arithmetic computation; (4) Multimodal Matching, represented by the COCO dataset (Le et al., 2023) for image-text alignment; (5) Code-based Tasks, including HumanEval-Exe (Chen et al., 2021) for execution simulation, Open-Critic (Vezora, 2024) for generation, and Code-Preference (Vezora, 2024) for summarization. These datasets are therefore intentionally positioned as a prerequisite stage to support safe and informed downstream application.
@@
Our Preprocessing. To support our decompositional evaluations, we curate those datasets to augment each instance with three additional aspects of information relevant to Tasks I–III (Section 3.2). Specifically, we begin by identifying and annotating the causal variables $(X,Z,M,Y)$ from the original data, questions, or descriptions. Using these annotations, we construct a DAG to represent the underlying causal structure of each data instance, which enrich original instances with causal and counterfactual structures. A running example in Figure 2.
Preprocessing Feasibility. Notably, all datasets are built upon “what-if” conditions or hypothetical scenarios (as outlined in Table 1) and intervention-style narratives, thus naturally supporting counterfactual interventions. For each instance, we parse and extract the intervened variables and, guided by the previously constructed DAG, annotate the corresponding counterfactual outcomes and construct a matched counterfactual graph. We provide the curated instances in Appendix A.
@@ 4 Experiment
We aim to empirically answer two research questions: RQ1: How well do LLMs perform when their counterfactual reasoning is decomposed into distinct reasoning tasks? RQ2: What auxiliary techniques can improve LLMs’ counterfactual reasoning? We defer the experimental settings into Appendix B.1 and additional results into Appendix B.2.
LLMs. We evaluate reasoning-centric and multimodal LLMs due to the nature of counterfactual reasoning tasks. Specifically, we leverage GPT-5, GPT-o4-mini-high, Qwen3-VL-235B-A22B-Thinking, Llama-4-Scout-17B, Llama-4-Maverick-17B-128E, Gemini2.5-Pro, DeepSeek-VL.
@@ 4.1 LLM Performance on Decompositional Tasks (RQ1)
Setting. We evaluate LLMs independently on each decompositional task. For each task, we explicitly provide the ground-truth outputs from the preceding tasks to isolate and measure LLMs’ capabilities specific to that task. For example, when assessing models’ ability to construct causal graphs, we supply the original inputs along with the ground-truth causal variables.
Table 2: LLMs’ performance in causal variable identification, we report means of F1 across all instances for each variable. Each value is scaled to 100%. The standard deviation is in Table 8.
Dataset GPT-5 GPT-o4 Qwen3 Llama4-S Llama4-M Gemini2.5 DeepSeek
$v_{1}$ $v_{2}$ $v_{1}$ $v_{2}$ $v_{1}$ $v_{2}$ $v_{1}$ $v_{2}$ $v_{1}$ $v_{2}$ $v_{1}$ $v_{2}$ $v_{1}$ $v_{2}$
$v_{1}=X$ (Exposure), $v_{2}=Z$ (Covariate)
CRASS 92.3 91.1 91.0 89.2 87.3 85.4 88.5 86.9 90.6 89.1 88.6 86.2 89.5 87.1
CLOMO 89.8 87.6 88.1 85.6 83.5 81.9 87.1 85.3 88.8 87.0 84.3 82.8 86.4 84.2
RNN-Topo 87.9 85.4 85.9 83.7 80.4 78.6 84.3 82.6 85.7 84.2 81.7 80.1 83.8 82.0
CVQA-Bool 79.4 76.2 79.8 76.5 72.3 70.5 77.9 75.8 79.1 76.9 68.5 66.9 70.7 68.3
CVQA-Count 74.7 72.3 74.3 72.9 68.9 67.2 73.6 71.9 74.4 72.5 65.8 63.7 67.4 65.1
COCO 72.8 70.2 73.2 71.1 67.2 65.4 72.5 70.8 73.6 71.7 62.6 60.9 65.9 63.2
Arithmetic 88.2 86.5 84.9 82.8 75.7 73.8 80.3 78.6 81.6 79.8 76.9 74.5 78.3 76.1
MalAlgoQA 84.1 81.3 81.5 78.9 72.9 70.6 79.5 77.1 80.6 78.2 73.5 71.2 75.9 73.4
HumanEval-Exe 69.3 66.9 71.4 69.2 63.7 61.9 67.8 65.7 68.9 66.7 59.6 57.3 62.1 59.8
Open-Critic 71.7 69.4 70.1 67.3 61.8 59.7 66.5 64.7 67.6 65.8 57.3 55.9 60.4 58.1
Code-Preference 49.6 68.4 80.2 69.0 72.9 60.5 73.6 61.9 75.2 63.4 68.4 66.5 61.2 59.3
$v_{1}=M$ (Mediator), $v_{2}=Y$ (Outcome)
CRASS 87.4 91.7 84.1 89.3 72.8 79.9 81.2 87.4 82.4 88.6 74.1 81.6 76.2 83.5
CLOMO 83.1 89.4 81.3 87.9 68.9 77.4 77.2 84.9 78.6 86.0 71.2 78.9 73.5 80.7
RNN-Topo 81.7 86.3 79.4 85.6 67.1 75.8 75.3 83.1 76.6 84.4 69.3 76.4 71.5 78.9
CVQA-Bool 73.5 79.3 73.9 80.1 59.7 66.9 71.8 78.3 72.9 79.4 57.3 63.2 60.5 68.4
CVQA-Count 69.6 75.2 70.1 76.2 56.8 63.6 68.9 74.8 70.2 75.9 54.7 60.4 57.9 65.1
COCO 67.3 73.4 68.0 74.4 54.2 61.8 66.2 72.1 67.4 73.5 52.8 58.3 55.7 62.6
Arithmetic 82.1 85.6 82.3 73.4 63.6 71.9 79.1 85.1 80.3 86.0 62.8 73.5 65.7 74.9
MalAlgoQA 79.2 83.4 76.8 80.4 61.5 69.3 76.6 81.2 77.8 82.4 60.2 70.5 63.8 72.3
HumanEval-Exe 63.2 67.4 66.0 70.3 51.9 59.7 61.9 66.3 63.0 67.5 49.7 57.0 53.6 60.5
Open-Critic 66.3 70.6 64.1 68.9 49.7 57.2 64.7 69.0 65.6 70.0 47.3 54.8 51.3 58.7
Code-Preference 65.9 75.3 66.3 77.0 50.4 78.6 64.5 79.3 65.7 80.4 48.2 76.1 52.5 59.8
Task I: Causal Variable Identification. Table 2 presents the performance of LLMs on identifying causal variables ($X,Z,M,Y$). We observe that model performance is strongly influenced by the modality of the dataset. Specifically, datasets involving more complex modalities (e.g., images, mathematical symbols, codes) tend to reduce LLM accuracy (e.g., <<0.7 F1 on Open-Critic), even when variables like $X,Z,Y$ are explicitly present in the context.
@@
Table 3: LLM performance (F1 mean) in reasoning the counterfactual mediator ($M^{\prime}$) and outcome ($Y^{\prime}$). Standard deviation is in Table 9.
Dataset GPT-5 GPT-o4 Qwen3 Llama4-S Llama4-M Gemini2.5 DeepSeek
$M^{\prime}$ $Y^{\prime}$ $M^{\prime}$ $Y^{\prime}$ $M^{\prime}$ $Y^{\prime}$ $M^{\prime}$ $Y^{\prime}$ $M^{\prime}$ $Y^{\prime}$ $M^{\prime}$ $Y^{\prime}$ $M^{\prime}$ $Y^{\prime}$
CRASS 92.1 88.0 90.5 86.2 80.5 73.9 70.1 63.5 84.9 79.5 81.7 75.2 82.9 77.1
CLOMO 90.2 85.3 88.7 83.9 77.8 71.6 67.2 60.9 82.9 77.2 79.3 72.8 80.5 74.3
RNN-Topo 88.9 83.4 87.9 81.6 75.6 69.4 65.2 58.7 80.5 75.0 77.1 70.6 78.3 72.0
CVQA-Bool 81.2 74.5 77.1 70.2 65.4 58.6 54.8 48.1 70.9 64.3 63.2 56.8 66.7 59.8
CVQA-Count 79.2 72.0 76.2 69.3 62.2 55.7 51.7 45.2 67.5 61.0 60.1 53.8 63.5 56.9
COCO 77.8 70.1 75.3 66.9 60.1 53.4 49.3 42.7 65.4 58.7 57.8 51.5 61.3 54.6
Arithmetic 87.8 82.7 85.8 80.9 69.8 63.2 59.1 52.4 76.3 70.6 72.1 65.4 74.0 67.5
MalAlgoQA 85.1 79.6 83.6 77.8 67.5 60.9 57.4 50.7 74.0 68.2 69.6 62.9 71.8 65.1
HumanEval-Exe 75.7 71.5 73.4 66.5 58.2 51.5 47.7 41.2 63.6 56.9 55.8 49.4 59.4 52.7
Open-Critic 75.3 69.4 73.8 67.5 56.0 49.4 45.8 39.2 61.5 54.7 53.7 47.3 57.2 50.6
Code-Preference 77.0 71.0 74.4 66.8 57.1 50.4 46.6 40.0 62.7 55.9 54.7 48.3 58.3 51.6